A firm has found a way of using first-degree price discrimination. Demand for its product is given by

P = 20 – 2Q

Marginal cost is constant and equal to $6.

i) With first-degree discrimination, what will be the profit-maximizing rate of output? How much economic profit will the firm earn?

ii) What will be the profit-maximizing rate of output if the firm does not discriminate and sets one price for all customers? How much economic profit will the firm earn in this case?



1
Expert's answer
2022-01-03T09:24:36-0500

i) "P=MC"

"20-2Q=6"

"14=2Q"

"Q=7"

"Profit = TR-TC"

"Profit=P*Q-TC"

"Price = 6"

"Profit = 7*6-6*6"

Profit is $6

ii) It will be higher than $6